Calculate the benefit-cost ratio, IRR, Annual Percentage Rate and Payback Period for the following project. Consider an inicial investment of 3,000,000 dollars:

Year Income Costs
1 200,000 140,000
2 200,000 140,000
3 200,000 140,000
4 200,000 140,000
5 200,000 140,000
6 200,000 140,000
7 200,000 140,000

An investor requires a 15% minimum acceptable rate of return. Is that possible?. Justify your answers.
Answer #1

А В C Е Present Value of Income Present Value of Cost Cost/1.15 year) 1 Year (Income/1.15Ayear) Income Cost 3,000,000 2 0 3,0

Benefit -Cost Ratio= Present Value of Benefit/Present Value of Cost= 832084/3582459=0.2322

The Net Cash flow and IRR calculation is given below:

А В C D Cash flow 1 Year (Income-Cost) Income Cost (3,000,000) 3,000,000 2 0 3 1 200,000 140,000 60,000 200,000 60,000 2 140,

В C Cash flow 1 Year (Income-Cost) Income Cost (3,000,000) 60,000 2 0 3,000,000 200,000 140,000 1 4 2 200,000 140,000 60,000

Hence, IRR is -33.766%

To calculate annual rate: PV=-3000000, nper=7, PMT=60000, rate=?

G Н J К (3,000,000) PV 7 nper РMT 60000 =RATE(H2,H3,H1,0,0)| rate RATE (nper, pmt, pv, [fv], [type], [guess])

G Н (3,000,000) PV 7 nper РMT 60000 -33.77% rate

Hence, Annual rate is -33.77%

As, the initial investment is $3 million and total net profit we get from this project is 60k*7=$420K. Hence,we do not recover our money in the investment.

As, the investor require 15% accpetable rate of return and we get -33.77% return, we should not invest in this project.